OpenStudy (anonymous):

Tonya goes grocery shopping and wants to buy some yogurt. She sees that a 32 ounce container of yogurt costs $2.69 and an 8 ounce container of yogurt costs $0.75. She decides to buy the 8 ounce yogurt, since it is so much cheaper. What mistake did she make here? Make sure to explain what Tonya did wrong and how she can find out which container of yogurt is the better buy. Show and explain all of your work, and clearly identify which one is the better buy.

OpenStudy (triciaal):

Tonya did not compare the price for the same unit. case 1: 32 ozs cost $2.69 case 2 8 oz cost 0.75 so 32 oz would cost (0.75 x 4) = $3.00 the better buy is case 1 since it cost 2.69/32 = 0.08 per oz rather than case 2 where it cost 0.75/8 = 0.09 per oz For every 32 oz purchase with case 1 she would save 3.00-2.69 = 0.31

OpenStudy (anonymous):

@triciaal Thank You! :)

OpenStudy (triciaal):

welcome